OpenStudy (anonymous):

find the inverse of the following function and write it in the form y=f⁻¹(x) f(x)=8-7x

OpenStudy (anonymous):

whenever looking for inverses, just switch the x and y positions then solve for y... in your problem, let f(x)=y=8-7x or just y = 8 - 7x.... switching the x and y will give you: x = 8 - 7y from here, just solve for y and that will give you the inverse of f, or \(\large f^{-1}(x) \)

OpenStudy (anonymous):

yeah i did that but it said it was wrong, they want the first firm in f⁻¹

OpenStudy (anonymous):

form*